OpenStudy (anonymous):

Pls explain how you got it so I can learn!! Factor this polynomial: 12x + 8 3(4x+2) 2(4x+3) 4(3x+2) 4(2x+3)

OpenStudy (anonymous):

just take 4 common

OpenStudy (anonymous):

12=4*3 and 8=4*2 =4*3*x+4*2

OpenStudy (anonymous):

after takin 4 out 4(3x+2)

OpenStudy (anonymous):

Ow... wow i so dumb but thank you very much ^-^ it help me a lot